OpenStudy (anonymous):

Find tanx=3cosx for all answers from [0,2pi)

OpenStudy (welshfella):

try writing tan x = sin x / cos x

OpenStudy (anonymous):

Welcome!

OpenStudy (welshfella):

sin x / cos x = 3 cos x 3 cos ^2 x = sin x now use the identity cos^x = 1 - sin^2x - substitute for cos^2 x to get an equation only in sin x

OpenStudy (anonymous):

I got the equation down to 3sin^2(x)+sin(x)-3=0 and used the quadratic equation. The result was( -1+- sqrt(37))/6. Does that seem correct?

OpenStudy (welshfella):

so we get 3( 1 - sin^2 x 0 = sin x 3 - 3sin^2 x = sin x 3sin^2 x + sin x - 3 = 0 solve this

OpenStudy (welshfella):

your equation was correct

OpenStudy (anonymous):

Now that I have this answer, how would I put it in terms of pi/radian form?

OpenStudy (welshfella):

yes that solution for sin x looks good

OpenStudy (welshfella):

well work the values out on a calculator first they come to 0.8471 and - 1.18 i would just give the answers as radians in decimal form note that 2pi = 6.283

OpenStudy (anonymous):

Thank you very much, this has been a huge help!

OpenStudy (welshfella):

yw bro